OpenStudy (anonymous):

find the critical numbers of, f(x)=8secx+4tanx, 0

OpenStudy (anonymous):

thanks guys im really freaking out abouit this..lol

OpenStudy (anonymous):

sec x = 1/cos x tan x = sin x/cos x f(x) = 8/cos x + 4sin x/cos x f(x) = (8+4sin x)/cos x We'll determine the derivative: f'(x) = [4(cos x)^2+ 8sin x + 4(sin x)^2]/(cos x)^2 But (cos x)^2 + (sin x)^2=1 f'(x) = (4+8sinx)/(cos x)^2 To find critical numbers f'(x) = 0 (4+8sinx)/(cos x)^2 = 0 4+8sin x=0 1+2sin x=0 2sin x=-1 sin x=-1/2 x=7pi/6 and x=11pi/6